OpenStudy (anonymous):

evaluate each determinant | 5 7| | 2 3 |

OpenStudy (ybarrap):

5*3 - 7*2 = 15-14 = 1

OpenStudy (anonymous):

thank you - no idea what that meant. so you just cross multiply and then subract what answers you get?

OpenStudy (ybarrap):

yes.

OpenStudy (ybarrap):

for 3x3 matrix and higher, little more complex, but not too hard

OpenStudy (anonymous):

thank you. SO much!

OpenStudy (ybarrap):

np

OpenStudy (anonymous):

I need a reason to use this in real life - when do you think you would?